OpenStudy (anonymous):

how to convert to higher terms 8/9= how do you get the answer

OpenStudy (anonymous):

the answer is 96/108

OpenStudy (anonymous):

how get that answer

OpenStudy (anonymous):

To get the answer you first need to realize that for example 3x/3y = (3/3)*(x/y) = x/y. So basically, what number can you multiply the 8 and 9 in 8/9 by to get 96/?. So you divide 96 by 8 to get 12. So that means you multiple the numerator and denominator by 12 to get (12/12)*(8/9) = 96/108.

OpenStudy (anonymous):

Thanks im currently in school and math is putting a work on me

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

I'm glad to help :)

OpenStudy (anonymous):

well every week ill be posting problems if you can help me with it

OpenStudy (anonymous):

can you also tell me about reducing to the lowest terms like 96/108

OpenStudy (anonymous):

Oh OK I will be around if you need me to help Oh and just go add me as a fan so you can send me messages if you need me.

OpenStudy (anonymous):

OK well you just convert 12 3/4 to improper fraction so 12 3/4 as an improper fraction is 51/4.